OpenStudy (anonymous):

When two atoms of 2 H (deuterium) are fused to form one atom of 4 He (helium), the total energy evolved is 3.83 × 10 -12 joules. What is the total change in mass (in kilograms) for this reaction?

OpenStudy (anonymous):

Please just give me the equation I need to use with the values plugged in, that's all I need.

OpenStudy (anonymous):

@Abhisar

OpenStudy (jfraser):

use \(E = mc^2\) Rearrange to solve for m, then plug in the energy and speed of light

OpenStudy (anonymous):

m = E/c^2 m = 3.83 x 10^-12/c^2 how do I find c^2?

OpenStudy (jfraser):

c is the speed of light

OpenStudy (jfraser):

square it

OpenStudy (anonymous):

8.98755179 × 10^16 m^2/s^2

OpenStudy (anonymous):

things are starting to come together...lmao m = E/c^2 m = 3.83 x 10^-12/8.98755179 × 10^16 m^2/s^2

OpenStudy (jfraser):

\( 9*10^{16} \frac{m^2}{s^2}\) would be enough

OpenStudy (jfraser):

but yes

OpenStudy (anonymous):

4.3 x 10^-29

OpenStudy (anonymous):

Thank you! I now understand it completely! :D :happydance:

OpenStudy (jfraser):

the last 4 questions you've asked have all been about \(E = mc^2\) just moving the pieces around

OpenStudy (jfraser):

there's really only 2 things to find: E or m, since c is always the speed of light

OpenStudy (anonymous):

I see that now, yeah. Now all the pieces are clicking. :D